The length of a rectangle is three feet shorter than twice its width. If the are of a rectangle is 170ft^2, what is the length of the rectangle?
Answer by mananth(16946)

The length of a rectangle is three feet shorter than twice its width. If the are of a rectangle is 170ft^2, what is the length of the rectangle?
...
let width be x feet
length = 2x-3
Area = 170 ft^2
...
Area = L*W
x(2x-3)=170
2x^2-3x=170
2x^2-3x-170 =0
2x^2-20x+17x-170=0
2x(x-10)+17(x-10)=0
(x-10)(2x+17)=0
Take positive value x= 10 feetthe width
Length = 20-3 = 17 feet
...
m.ananth@hotmail.ca